The two mathematical models that apply to the standing waves on strings system are the relationship between wave speed, wavelength, and frequency

and equation 1 in the handout relating the wave speed to the tension and linear density in the string. If you create a graph of wavelength as a function of string tension, you should expect a trend line close to _____.

a. exponential function

b. a power function of order 2

c.  power function of order 0.5
Answer #1

V =   √ T/μ \lambda =V/f = (√ T/μ )/f so \lambda is a function of square root that power function 0.5 option c